OpenStudy (anonymous):

identify a pattern arithmetic 9 19 29 39 49 ( ) ( ) ( ) wat gose in the blank

OpenStudy (anonymous):

59 69 79

OpenStudy (anonymous):

thxs

jimthompson5910 (jim_thompson5910):

You're adding 10 each time since 19-9 = 10, 29-19 = 10, etc... So the next term is 49+10 = 59 The term after that is 59+10 = 69 and the term after that is 69+10 = 79